OpenStudy (anonymous):

three of four numbers have sum of 22. If the average of the four numbers is 8, what is the fourth number

OpenStudy (anonymous):

Firstly you must know that: \[Average = \frac{Sum \; Of \; Numbers}{Total \;Numbers}\]

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!!!}\)

OpenStudy (anonymous):

If I have four numbers, let us say, x, y, z, h, then: Sum Of Numbers = x + y + z + h And you are just given that three of four numbers have sum of 22.. So, I am assuming that sum of x, y and z is 22.. So: Average = (x + y + z) + h = 22 + h And as you have 4 numbers(x,y,z,h), So Total Numbers = 4..

OpenStudy (anonymous):

Also, you have told that Average = 8.. So, just use that formula: \[8 = \frac{22 + h}{4} \implies 8 \times 4 = 22 + h \implies \color{blue}{32 - 22 = h} \implies \color{green}{h = 10}\] So, the fourth number is 10..

OpenStudy (anonymous):

And yes, lastly: \[\huge \color{green}{\textsf{Welcome To Openstudy...}}\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

While I do, you must respond or communicate, this way you will get things easily and more.. :)